OpenStudy (anonymous):

For exponential funcitons of the form f(x) = a^x which of the following expressions is equal to f (3) x f(4) answer choices: A: f(3^4) B: f(3 x 4) C: f(3 + 4)

OpenStudy (freckles):

well try to plug in let's find out f(3)=a^3 f(4)=a^4 if you multiply those what do you get?

OpenStudy (anonymous):

idk wats the answer?

OpenStudy (anonymous):

do you know the properties of exponents? Like: \[a^x \cdot a^y = a^{x+y}\]

OpenStudy (anonymous):

whats the answer a or b or c?

OpenStudy (anonymous):

i'm not here to give out answers... i'm here to help people learn. doesn't seem like you belong to the latter group. try reading the code of conduct (below, left). good luck

OpenStudy (freckles):

I'm not interested in giving just the answer this question isn't that bad if you try

OpenStudy (anonymous):

I am so sorry ill try

OpenStudy (anonymous):

im just in a hurry

OpenStudy (anonymous):

the finish line of life is death... take your time and enjoy what time you have

OpenStudy (anonymous):

what's f(3) = ?

OpenStudy (anonymous):

what does f(4)= ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

what expression do you get when you multiply them together?

OpenStudy (anonymous):

which property of exponents can you use to simply the expression?

OpenStudy (anonymous):

rewrite that expression using the function notation.

OpenStudy (anonymous):

hey I think the answer is A am I right!!

OpenStudy (anonymous):

which answer matches your function?

OpenStudy (anonymous):

done!

OpenStudy (anonymous):

i got it right?

OpenStudy (anonymous):

nvm got it wrong.

OpenStudy (anonymous):

i like alot of chicken niggets